OpenStudy (agentjamesbond007):

Related Rates: Two sides of a triangle are 4 m and 5 m in length and the angle between them is increasing at a rate of 0.06 rad/s. Find the rate at which the area of the triangle is increasing when the angle between the sides of fixed length is π/3.

OpenStudy (anonymous):

A=1/2(4)(5)(sin of the angle) Differentiate with repect to time, dA/dt = 10 x d(sine of the angle)/dt = 10 x (cosine of angle) x d(angle)/dt = 10 x 1/2 x 0.06

OpenStudy (agentjamesbond007):

Thank You! From where did you get the sin of the angle?